OpenStudy (anonymous):

What is the equation of the given line in standard form? Use integer coefficients.y = -6.9x + 5.1

OpenStudy (anonymous):

A. -69x + 10y = 51 B. 69x + 10y = 51 C. 69x + 10y = -51 D. -69x + 10y = -51

OpenStudy (aripotta):

ok. what form is it in now?

OpenStudy (anonymous):

@AriPotta what do you mean

OpenStudy (aripotta):

i want to make sure that you know the different forms for equations of lines. what form is y = -6.9x + 5.1 written in?

OpenStudy (anonymous):

i d know i dont even know what your asking

OpenStudy (aripotta):

there's standard form, slope-intercept form, and point-slope form.

OpenStudy (anonymous):

this is pont- slope form right

OpenStudy (aripotta):

no :/ point-slope form is y - y1 = m(x - x1)

OpenStudy (anonymous):

o haha im stupid

OpenStudy (aripotta):

ok, so it's obviously in slope-intercept form, which is y = mx + b

OpenStudy (aripotta):

do you know what standard form is?

OpenStudy (anonymous):

@AriPotta no

OpenStudy (aripotta):

standard form: Ax + By = C

OpenStudy (aripotta):

and A cannot be negative or a fraction

OpenStudy (aripotta):

so how can we get y = -6.9x + 5.1 into standard form?

OpenStudy (anonymous):

by adding 6.9 to each side?

OpenStudy (aripotta):

right, but you gotta include the x. so what do you get after adding 6.9x to both sides?

OpenStudy (anonymous):

12x

OpenStudy (aripotta):

12x? how'd you get 12x?

OpenStudy (anonymous):

i add both sides like you said

OpenStudy (aripotta):